### v1

# Draft Model Knows When to Stop: A Self-Verification Length Policy for Speculative Decoding

Abstract
Speculative Decoding (SD) has become an important technique in accelerating the inference speed of large language models. Conventional SD methods employ a fixed draft length, which ignores the token generation difficulty across tasks. Consequently, in this paper, we address such an issue and introduce SVIP - a difficulty-aware dynamic draft length policy for speculative decoding systems. Based on a theoretical lower bound of draft token acceptance rate and its inference-time approximation, SVIP adaptively determines the lengths of draft sequences based on the entropy of each draft token distribution. Experimental results on mainstream SD benchmarks and frameworks demonstrate the superior performance of SVIP, achieving up to 20% walltime speedup on SpecBench over baseline SD methods and 60% speedup on MT-Bench for long-form generation of up to 8K tokens. Moreover, SVIP is totally training-free and compatible with any existing SD methods that generate draft tokens autoregressively. Experimental results also show that SVIP yields consistent walltime improvement on top of GliDe & CaPE and EAGLE-2.
Draft Model Knows When to Stop:
A Self-Verification Length Policy for Speculative Decoding
Ziyin Zhang1,2thanks: Work done during their internship at Tencent AI Lab. Jiahao Xu2 Tian Liang2 Xingyu Chen1,2∗ Zhiwei He1,2∗ Rui Wang1thanks: Corresponding author.Zhaopeng Tu2 1Shanghai Jiao Tong University 2Tencent AI Lab 1{daenerystargaryen,galaxychen,zwhe.cs,wangrui12}@sjtu.edu.cn 2{jettexu,ttianliang,zptu}@tencent.com
1 Introduction
Speculative decoding (Leviathan et al., 2023; Chen et al., 2023) is a novel technique that markedly enhances the generation wall-time of large language models (LLMs). This approach employs a small and efficient amateur model to draft sequences, while concurrently utilizing a larger and more powerful expert model to verify the drafts. By avoiding the autoregressive generation of each token through the target LLM, speculative decoding achieves improved efficiency while preserving the quality of the output.
Figure 1: The “difficulty” of tokens varies in a sequence, resulting in different numbers of accepted draft tokens at different positions.
Many variants of speculative decoding have been proposed. A line of work focuses on developing a stronger and faster draft model (Li et al., 2024b; Elhoushi et al., 2024; Du et al., 2024). Another line of work contributes to maximizing the acceptance probability of draft tokens (Sun et al., 2023; Li et al., 2024a; Lu et al., 2024). In general, they all tend to maximize the alignment between the draft and target model to further maximize the system acceptance rate.
Though successful, most of these works limit their settings to a fixed draft length, where the draft model always generates a fixed number of tokens (e.g. 4 or 5) in each iteration. Such a setting ignores the fact that some tokens - such as stop words or civilities - in the generation may be easy for the draft model to predict, while others - such as knowledge-intensive or reasoning-intensive tokens - can be much harder, as shown in Figure 1.
Figure 2: The correlation between draft model entropy and draft token acceptance probability (top) and lengths of accepted draft seqeunces (bottom).
To address this issue, in this work we introduce SVIP - Self-VerIfication length Policy, a simple, plug-and-play dynamic draft length policy for speculative decoding systems, which enhances the wall time speedup of these systems by adaptively allowing for longer draft sequences for “simple” tokens (top of Figure 1) and terminating the drafting process early upon encountering “hard” tokens (bottom of Figure 1).
Specifically, we first analyze the acceptance rate of speculative decoding systems and derive a lower bound based on the system’s entropy information. Further empirical analysis suggests that such a bound can be approximated by the entropy of the draft model only, which is naturally available in the drafting process of any auto-regressive draft models. Consequently, we develop SVIP  which controls the length of draft sequences dynamically by determining whether to continue drafting or start verification upon the generation of each draft token.
With extensive experiments across multiple model sizes and generation lengths, we demonstrate the superior performance of SVIP. It yields more than 20% of improvements over vanilla speculative decoding for Qwen2.5 14B and LLaMA-3 70B on SpecBench (Xia et al., 2024), and more than 60% of improvements for Pythia 6.9B on MT-Bench (Zheng et al., 2023) when generating long-form responses of up to 8K tokens.
Moreover, since our method is lightweight and training-free, it is extremely flexible and can be adapted to any speculative decoding system with an auto-regressive draft model. As examples, we apply SVIP on top of two state-of-the-art speculative decoding methods: GliDe with a CaPE (Du et al., 2024) and EAGLE-2 (Li et al., 2024a), and confirm that it brings consistent improvements. Our code is available at https://github.com/Geralt-Targaryen/SVIP.
In summary, our contributions are threefold:
  1. 1.
    We derive a low bound of speculative decoding systems, where the acceptance rate of the draft model could be modeled by its entropy only.
  2. 2.
    Based on this lower bound, we further develop an entropy-based dynamic draft length policy for speculative decoding systems, which is extremely flexible and can be adapted to any auto-regressive draft model.
  3. 3.
    Experimental results demonstrate the superior performance of SVIP over baseline methods, with up to 20% average speedup on SpecBench, 60% in long-form generation, and consistent improvement over state-of-the-art speculative decoding frameworks such as GliDe & CaPE and EAGLE-2.
2 Method
Figure 3: Distribution histograms of the entropy ratio $H_{q,p}/H_{q}$. For most tokens, this ratio falls into a narrow range, indicating that the cross entropy $H_{q,p}$ can be approximated by a constant multiplication of the draft entropy $H_{q}$.
The overall objective of SVIP is to dynamically adapt draft length on-the-fly, stopping early if the current draft token’s acceptance probability is low and otherwise continuing drafting. To introduce our method, we first provide the background on speculative decoding in Section 2.1. Then, based on the key observation that the acceptance probability of a draft token depends on the target model confidence - which is unavailable in the drafting phase, we derive a theoretical lower bound for the acceptance rate based on the draft model’s entropy and the cross-entropy between the target and draft models in Section 2.2. Finally, based on empirical analysis of the target and draft distributions of more than 100K tokens across three model families, we approximate this lower bound using only the entropy information of the draft model in Section 2.3, making it viable in actual inference.
2.1 Preliminaries on Speculative Decoding
Suppose we have two LLMs $p$ and $q$, where $p$ is the larger (target) model, and $q$ is the smaller (draft) model. Given an input sequence $x_{\leqslant t}$ of length $t$, and a draft length $\gamma$, the draft model first samples $\gamma$ tokens $x_{t+1},\cdots,x_{t+\gamma}$ autoregressive, which are verified by the target models in parallel to acquire the confidences $p(x_{t+1}),\cdots,p(x_{t+\gamma})$ [^1] The short hand $p(x_{n})$ is used to denote the conditional probability $p(x_{n}|x_{<n})$ when there is no ambiguity. Throughout the work we use subscripts to indicate token indices in a sequence (e.g. $x_{n}$ for $n$-th token), and superscripts to indicate element indices in a vector (e.g. $x_{n}^{i}$ for $i$-th element in $x_{n}$)..
[…]
The complete algorithms for speculative decoding are given in Appendix A.
Figure 4: Comparison between the actual acceptance probability from Equation (2), the acceptance probability lower bound from Equation (5), and the estimated lower bound after approximating the cross entropy $H_{q,p}$ with a constant multiplication of $H_{q}$. Each position on the x-axis corresponds to a token, which has been sorted according to the actual acceptance probability.
2.2 Theoretical Lower Bound of Acceptance Rate
From Section 2.1, it’s easy to derive that given an input sequence $x_{<t}$ and a draft token $x_{t}$, its acceptance probability is $\min\left(1,\;\frac{p(x_{t})}{q(x_{t})}\right)$. Let $\beta$ denote the expected acceptance probability over the distribution of $x_{t}$, and it follows that
$\beta$ $=\sum_{x}q(x)\cdot\min\left(1,\frac{p(x)}{q(x)}\right)$ (1)
$=\sum_{x}\min\left(p(x),q(x)\right).$ (2)
Chen et al. (2023) has proven that $\beta$ is related to the total variational distance (TVD) between $p$ and $q$:
$\beta=1-\mathrm{TVD}(p,q).$ (3)
According to Pinsker’s inequality - which relates TVD to Kullback-Leibler divergence - we then have
$\beta$ $\geqslant 1-\sqrt{\frac{1}{2}\mathrm{KL}(q||p)}$ (4)
$=1-\sqrt{\frac{1}{2}\sum_{x}q(x)\log\frac{q(x)}{p(x)}}$
$=1-\sqrt{\frac{1}{2}H_{q,p}-\frac{1}{2}H_{q}},$ (5)
where $H_{q,p}$ is the cross entropy between $q$ and $p$, and $H_{q}$ is the entropy of $q$.
Equation (5) provides a theoretical lower bound for the acceptance rate of a draft token using the entropy information of the speculative decoding system [^2] An alternative way to construct the lower bound is discussed in Appendix B.. To provide an intuitive motivation for using entropy to construct the lower bound, we plot the relation between draft model entropy and draft token acceptance probability as well as accepted draft seqeunce lengths in Figure 2, which are collected from more than 1M tokens generated by three different target models with temperature set to 1 and maximum draft length set to 40. The first subfigure indicates a strong negative correlation between draft model entropy and draft token acceptance probability, while the second subfigure shows that when the entropy is low, dozens of consecutive draft tokens could be accepted, which highlights the drawback of setting draft length to a constant, small value, as is the common practice in speculative decoding literature. Another version of Figure 2 in the greedy setting is given in Appendix D.
2.3 Empirical Estimation of Acceptance Rate
So far in this section, we have been assuming access to the target distribution $p(x_{t})$ of the next token, which is unavailable in the drafting phase at inference time. Thus, to apply Equation (5) for actual acceptance rate estimation, we must approximate the cross entropy $H_{q,p}$ with information from only the draft model.
Algorithm 1 SVIP
1:target model $p$, draft model $q$, input sequence $x_{\leqslant t}$, maximum length $T$, threshold $h$
2:Initialize $n\leftarrow t$
3:while $n<T$ do
4:     $j=0$
5:     while True do
6:         Sample $x_{n+j}\sim q(x|x_{<{n+j}})$
7:         $j\leftarrow j+1$
8:         if $\sqrt{H(q_{x|x_{<n+j}})}>h$ then
9:              Exit while loop
10:         end if
11:     end while
12:     $\gamma\leftarrow j$
13:     Compute $p(x|x_{<{n+j}}),\;j=1,\cdots,\gamma+1$ in parallel
14:     $\tilde{n}\leftarrow n$
15:     for $j=1$ to $\gamma$ do
16:         if Verify$\left(p_{x|x_{<n+j}},\,q_{x|x_{<n+j}},x_{n+j}\right)$ then
17:              $\tilde{n}\leftarrow\tilde{n}+1$
18:         else
19:              $x_{n+j}\leftarrow\text{Correct}\left(p_{x|x_{<n+j}},\,q_{x|x_{<n+j}}\right)$
20:              Exit for loop
21:         end if
22:     end for
23:     if $\tilde{n}==n+\gamma$ then
24:         Sample $x_{n+\gamma+1}$ from $p(x|x_{\leqslant n+\gamma})$
25:     end if
26:     $n\leftarrow\tilde{n}+1$
27:end while
28:$x_{\leqslant n}$
To tackle this issue, we first plot the relationship between $H_{q,p}$ and $H_{q}$ in Figure 3. For all three model families, the entropy ratio $H_{q,p}/H_{q}$ are concentrated in a narrow range between 1.0 and 1.3. Thus, we choose to approximate $H_{q,p}$ with a constant multiplication of $H_{q}$. Plugging it into Equation (5), we now have
$\beta$ $\geqslant 1-\sqrt{cH_{q}},$ (6)
where $c$ is a constant controlling the approximation ratio between $H_{q,p}$ and $H_{q}$. In Figure 4, we visualize the values derived from Equation (2), (5), and (6).
With Equation (6) providing a way to estimate acceptance probability using only the draft model’s entropy, we can now adapt the draft length on-the-fly. After generating each draft token, we compute the estimated acceptance probability lower bound, and stop the draft process if it’s lower than a certain threshold $h$. We note that since both $c$ and $h$ are constant hyperparameters, we can remove $\sqrt{c}$ from Equation 6 and absorb it into the threshold $h$.
We formalize SVIP in Algorithm 1. The details of the methods Verify and Correct are given in Appendix A, for which different versions are available for sampling (Algorithm 2, 4) and greedy decoding (Algorithm 3, 5).
3 Experiments
3.1 Experiments on SpecBench
3.1.1 Settings
Methods MT-Bench Trans. Sum. QA Math RAG Avg.
Pythia (6.9B, ${}_{\text{(6.9B, 160M)}}$ Const. 1.45 1.47 1.24 1.43 1.52 1.42 1.42
Heuristics 1.51 1.58 1.34 1.58 1.64 1.51 1.53
SVIP 1.63 1.62 1.45 1.67 1.72 1.66 1.63(+14.8%)
Qwen2.5 (14B, ${}_{\text{(14B, 0.5B)}}$ Const. 1.08 0.87 1.11 0.92 1.43 0.99 1.07
Heuristics 1.10 0.91 1.10 0.92 1.34 1.03 1.07
SVIP 1.33 1.12 1.37 1.14 1.57 1.23 1.29(+20.6%)
LLaMA-3 (70B, ${}_{\text{(70B, 8B)}}$ Const. 2.04 2.48 2.56 2.34 2.32 2.28 2.34
Heuristics 2.30 3.13 3.33 2.61 2.52 2.63 2.76
SVIP 2.31 3.04 3.48 2.63 2.89 2.59 2.83(+20.9%)
Table 1: Speedup over target-model-only autoregressive decoding on SpecBench.
We validate the effectiveness of SVIP on Spec-Bench (Xia et al., 2024) using three distinct target models: Pythia 6.9B (Biderman et al., 2023), Qwen2.5 14B (Yang et al., 2024), and LLaMA-3 70B (Dubey et al., 2024), with Pythia-160M, Qwen2.5 0.5B, and LLaMA-3 8B as the draft models respectively.
As baselines, we consider two simple policies for draft length: 1) a constant draft length of 5, which is commonly used in the literature, and 2) the heuristics implemented in Hugging Face Transformers library (Wolf et al., 2019), where the draft length for the next draft iteration is increased by 2 if all draft tokens in the current iteration are accepted, and otherwise decreased by 1.
We set the sampling temperature to 0 on SpecBench (the alternatives are discussed in Appendix C). For each model, the entropy threshold $t$ in SVIP is chosen from $\{0.2,0.3,0.4,0.5\}$ based on performance on 8 samples held out from MT-Bench (Zheng et al., 2023), which are 0.4 for Pythia, and 0.3 for Qwen2.5 and LLaMA-3. All experiments with Pythia and Qwen are conducted on a single 40GB A100, while experiments with LLaMA are conducted on 5 40GB A100s. To mitigate the impact of system performance varitations, we repeat all experiments with Pythia and Qwen for three times (using different random seeds when they are used) and report the average speedup over target-model-only autoregressive decoding. Also, since the memory consumption of verifying $n$ draft tokens is quadratic in $n$, we limit the maximum draft length to 40 in both heuristics and SVIP scenarios, beyond which we start to encounter out-of-memory issues.
3.1.2 Results
The results on SpecBench are shown in Table 1. Compared with the constant approach, SVIP yields an average speedup of 15% for Pythia and 20% for Qwen and LLaMA. Compared with the heuristic approach, SVIP also gives a consistent improvement on all domains for Pythia and Qwen, and outperforms the latter on 4 out of 6 domains for LLaMA.
Figure 5: The average generated draft length, accepted draft length, and acceptance rate of Qwen2.5 (top) and LLaMA-3 (bottom) on SpecBench. Compared with the two baselines, SVIP leads to a shorter draft length and a much higher acceptance rate.
In Figure 5, we plot the average draft length and accepted draft length of Qwen and LLaMA (the results for Pythia are similar, and are given in Appendix C). From the figure, we observe that by terminating the draft process when the draft model entropy is high, SVIP leads to shorter draft lengths and a much higher acceptance rate (close to or more than 80% on all domains for Qwen, and more than 90% for LLaMA). Notably, the acceptance rate of LLaMA-3 even reaches 99% on the summarization domain, which contributes to the highest speedup (3.48) in Table 1.
3.2 Long-form Generation
Most existing works on speculative decoding (Chen et al., 2023; Du et al., 2024) limit their experiments to generating short sequences of 128 tokens. To verify the wide applicability of SVIP, we also conduct experiments on long-form generation with up to 8K context. For this purpose, we use MT-Bench (Zheng et al., 2023) as the dataset and set the sampling temperature to 1, as we found that when using greedy decoding in long-form generation, both the draft and the target models are prone to repeat themselves, resulting in very low information entropy (see Appendix D for details). Other settings follow Section 3.1.
Methods Generation Length
128 256 512 1K 2K 4K 6K 8K
Pythia (6.9B, ${}_{\text{(6.9B, 160M)}}$ Const. 0.68 0.69 0.69 0.70 0.72 0.90 0.89 0.87
Heuristics 0.88 0.88 0.88 0.88 0.90 1.21 1.25 1.23
SVIP 1.07 1.08 1.08 1.07 1.08 1.43 1.44 1.41
Qwen2.5 (14B, ${}_{\text{(14B, 0.5B)}}$ Const. 0.98 0.97 0.95 0.96 0.98 1.00 1.02 1.04
Heuristics 1.01 0.99 0.98 1.00 1.02 1.03 1.04 1.06
SVIP 1.29 1.29 1.30 1.31 1.32 1.33 1.33 1.35
LLaMA-3 (70B, ${}_{\text{(70B, 8B)}}$ Const. 1.74 1.74 1.77 1.78 1.82 1.90 1.93 1.94
Heuristics 1.53 1.56 1.61 1.63 1.68 1.77 1.83 1.84
SVIP 1.69 1.72 1.75 1.78 1.86 1.96 2.01 2.02
Table 2: Speedup on MT-Bench with different generation length.
The results are given in Table 2. Interestingly, we find that in the sampling setting, the two baseline methods (constant and heuristics) perform even worse than target-model-only auto-regressive decoding for Pythia and Qwen, while SVIP consistently yields a positive speedup. For LLaMA, while constant draft length performs better with contexts shorter than 1K tokens, SVIP exceeds it for generating longer sequences. Another observation is that the speedup ratio of all three methods generally increases with the context length, which could be possibly attributed to the longer contexts giving the draft model more information, aligning it at test time to the target model. However, we note that the absolute values of token throughput (which are not discussed in this paper, since they heavily depend on the underlying machines) stay at the same level from 1K to 8K context for Qwen2.5 and LLaMA-3 when using speculative decoding, and decrease slowly when not using speculative decoding. For Pythia, which does not use any optimized attention such as MQA (Shazeer, 2019) or GQA (Shazeer, 2019), the throughput decreases notably with context length in both cases - with or without specualtive decoding.
3.3 Applying SVIP to Other Draft Methods
In Section 3.1 and 3.2, we evaluated SVIP on vanilla speculative decoding, where a standard pretrained Transformer decoder model from the target model’s family is used as the draft model. However, in the past years many works on speculative decoding have proposed other stronger or more efficient draft models (Cai et al., 2024; Du et al., 2024; Li et al., 2024b). Since most of these works assume a constant draft length, SVIP is orthogonal to them and can be applied on top of them without any additional training.
Methods MT-Bench Code Finance GSM Spider Avg.
7B GliDe 1.95 2.04 1.91 1.98 1.69 1.95
   +SVIP 2.00 2.12 2.03 2.01 1.63 2.02
GliDe + CaPE 2.36 2.57 2.29 2.51 1.97 2.40
   +SVIP 2.56 2.65 2.49 2.54 2.08 2.52
13B GliDe 2.22 2.41 2.15 2.31 1.85 2.24
   +SVIP 2.31 2.43 2.17 2.35 1.85 2.28
GliDe + CaPE 2.73 2.86 2.66 2.80 2.24 2.73
   +SVIP 2.72 2.93 2.66 2.85 2.27 2.76
33B GliDe 2.12 2.25 2.09 2.29 1.99 2.18
   +SVIP 2.29 2.40 2.20 2.42 2.03 2.30
GliDe + CaPE 2.08 1.98 2.10 2.13 1.76 2.03
   +SVIP 2.13 2.02 2.15 2.16 1.82 2.08
Table 3: Speedup comparison with GliDe & CaPE, using Vicuna as the base model.
Specifically, we consider GliDe with a CaPE (Du et al., 2024), where the draft model - named GliDe - is a small transformer decoder with cross-attention to the target model’s hidden representations, while CaPE is a complementary tree expansion method to increase the acceptance rate of the draft token at each position. Following the settings of Du et al. (2024), we use Vicuna 7B, 13B, and 33B (Chiang et al., 2023) as the base models (for which the draft models are publicly available) and set the sampling temperature to 0. However, we distribute the 33B model across two GPUs due to memory constraint. Similar to the previous experiments, we set the threshold $t$ to 0.5 based on pilot experiments on 8 samples held out from MT-Bench using GliDe only (without CaPE).
We also apply SVIP to EAGLE-2 (Li et al., 2024a), the state-of-the-art speculative decoding system which utilizes the target model’s language modeling head on top of the draft model’s features to predict the next draft token, and dynamically constructs a draft tree at each draft position. Following Li et al. (2024a), we use LLaMA-2 7B, 13B (Touvron et al., 2023) and Vicuna 7B, 13B (Chiang et al., 2023) as the base models, and set the sampling temperature to 1. As Brown et al. (2024) suggest, adding a conditional clause (that decides whether or not to stop drafting) after the generation of each draft token in EAGLE-2 may interrupt the otherwise input-independent control flow of EALGE-2, reducing the effects of low-level interpreter and system optimizations. Thus, in EAGLE-2 we calculate the draft entropy and decide whether or not to stop drafting after every two draft tokens instead of after every draft token, as used in previous experiments.
Methods MT-Bench H-Eval GSM8K Alpaca CNN/DM QA Avg.
LLaMA-2 7B EAGLE-2 3.10 3.61 3.15 3.10 2.76 2.84 3.10
   + SVIP 3.16 3.66 3.18 3.13 2.82 3.02 3.16
LLaMA-2 13B EAGLE-2 3.38 4.12 3.41 3.25 3.01 2.98 3.41
   + SVIP 3.41 4.09 3.45 3.34 3.05 3.17 3.46
Vicuna 7B EAGLE-2 2.66 2.84 2.77 2.48 2.31 2.13 2.58
   + SVIP 2.84 2.97 2.75 2.66 2.42 2.30 2.73
Vicuna 13B EAGLE-2 2.85 3.31 2.93 2.74 2.48 2.30 2.83
   + SVIP 2.94 3.49 3.19 2.89 2.60 2.53 2.98
Table 4: Speedup comparison with EAGLE-2, using LLaMA-2-Chat and Vicuna as the base models.
The results for these two methods are presented in Table 3 and 4, respectively. For GliDe, SVIP yields a consistent speedup both with and without CaPE, with a 5% improvement for the 7B and 33B models. For EAGLE-2, the speedup is also particularly notable for the Vicuna models, with 5% improvement for both the 7B and 13B models.
4 Related Work
Since Leviathan et al. (2023) and Chen et al. (2023) introduced speculative decoding into large language models, numerous works have followed their tracks in pursuit of more efficient LLM inference. We broadly categorize these works into three types: better draft models, draft tree expansion, and draft length control, which are orthogonal to each other. A more comprehensive review of speculative decoding is provided by Xia et al. (2024).
Better draft models.
As Xia et al. (2024) suggest, draft models in speculative decoding can be either based on self-drafting or based on an independent draft model. For the first type, one may use a quantized (Zhao et al., 2024), early-exiting (Elhoushi et al., 2024), or forward-padded (Monea et al., 2023) version of the target model to produce draft tokens, while the second type is represented by the vanilla speculative decoding (Leviathan et al., 2023). Some works also take the best of both worlds and introduce extra layers on top of the target model’s hidden representations to construct draft models, represented by EAGLE (Li et al., 2024b), GliDe (Du et al., 2024), and Medusa (Cai et al., 2024).
Draft tree expansion.
Given a draft model, one may verify multiple draft tokens for the same position in parallel to increase the probability of finding an accepted draft token, and we use “draft tree expansion” as an umbrella term for such techniques. Li et al. (2024a) introduce EAGLE-2, which reranks draft tokens in EAGLE’s draft tree to select tokens with the highest confidence for verification. Similarly, CaPE (Du et al., 2024) improves GliDe by expanding the token set chosen for verification at each position based on top-1 confidence. Other works have also addressed the problem of multi-draft verification from a theoretic perspective (Sun et al., 2023; Yin et al., 2024).
Draft length control.
Works in this category are few, but most relevant to ours. Liu et al. (2024) introduce PEARL, which lets the target model perform verification in parallel to draft generation, stopping the draft process when a mismatch is found. Huang et al. (2024) propose SpecDec++, which trains an acceptance prediction head on top of the draft model to predict the acceptance probability of the current draft token, stopping the draft round when the predicted acceptance probability falls below a constant threshold. Brown et al. (2024) propose Dynamic Depth Decoding on top of EAGLE-2, which uses the sum of all tokens’ confidences in one level of its draft tree as an indicator to predict whether or not to continue draft generation.
5 Conclusion
We propose SVIP, a flexible, training-free, and plug-and-play dynamic draft length policy for speculative decoding systems. Based on a theoretical lower bound of acceptance probability and its empirical approximation, SVIP determines whether to continue draft generation or to quit drafting based on the draft model’s entropy after the generation of each draft token. With extensive experiments spanning various base models, draft methods, test domains, and generation length, we validated the effectiveness of SVIP, sparking new insights on speculative decoding and more efficient large language models.
Appendix A The Complete Speculative Decoding Algorithms
In Algorithm 2 to 6, we present the complete algorithms of the vanilla speculative decoding in both the greedy decoding and the sampling scenarios. For the sampling scenario, the Verify and Correct methods in Algorithm 6 resolve to Algorithm 2 and 4. For greedy decoding, they resolve to Algorithm 3 and 5.
Algorithm 2 Verify (Sampling)
1:target distribution $p(x)$, draft distribution $q(x)$, draft token $x_{t}$
2:$accept\leftarrow$ False
3:$r\sim U[0,1]$
4:if $r<\frac{p(x_{t})}{q(x_{t})}$ then
5:     $accept\leftarrow$ True
6:end if
7:$accept$
Algorithm 3 Verify (Greedy)
1:target distribution $p(x)$, draft distribution $q(x)$, draft token $x_{t}$
2:$accept\leftarrow$ False
3:if $\arg\max p(x)==x_{t}$ then
4:     $accept\leftarrow$ True
5:end if
6:$accept$
Algorithm 4 Correct (Sampling)
1:target distribution $p(x)$, draft distribution $q(x)$
[…]
3:$\hat{x}$
Algorithm 5 Correct (Greedy)
1:target distribution $p(x)$, draft distribution $q(x)$
2:$\arg\max p(x)$
Algorithm 6 Speculative Decoding
1:target model $p$, draft model $q$, input sequence $x_{\leqslant t}$, maximum length $T$, draft length $\gamma$
2:Initialize $n\leftarrow t$
3:while $n<T$ do
4:     for $j=1$ to $\gamma$ do
5:         Sample $x_{n+j}\sim q(x|x_{<{n+j}})$
6:     end for
7:     Compute $p(x|x_{<{n+j}}),\;j=1,\cdots,\gamma+1$ in parallel
8:     $\tilde{n}\leftarrow n$
9:     for $j=1$ to $\gamma$ do
10:         if Verify$\left(p(x|x_{<n+j}),\,q(x|x_{<n+j}),x_{n+j}\right)$ then
11:              $\tilde{n}\leftarrow\tilde{n}+1$
12:         else
13:              $x_{n+j}\leftarrow\text{Correct}\left(p(x|x_{<n+j}),\,q(x|x_{<n+j})\right)$
14:              Exit for loop
15:         end if
16:     end for
17:     if $\tilde{n}==n+\gamma$ then
18:         $x_{n+\gamma+1}\sim p(x|x_{\leqslant n+\gamma})$
19:     end if
20:     $n\leftarrow\tilde{n}+1$
21:end while
22:$x_{\leqslant n}$
Appendix B Alternatives for Acceptance Rate Lower Bound Computation
In Section 2, we used Pinsker’s inequality to compute a lower bound for the expected acceptance probability:
$\beta$ $=\sum_{x}\min\left(p(x),q(x)\right)$ (7)
$\geqslant 1-\sqrt{\frac{1}{2}KL(q||p)}.$ (8)
Another way to compute the lower bound of acceptance probability can be derived from Bretagnolle-Huber inequality (Bretagnolle and Huber, 1978):
$\beta$ $\geqslant 1-\sqrt{1-e^{-KL(q||p)}}.$ (9)
Compared with the Pinsker’s bound, it’s trivial to see that this bound is guaranteed to be always larger than 0. However, in practice we find that the Pinsker’s bound is 11% tighter for Qwen2.5, 20% tighter for Pythia, and 43% tighter for LLaMA-3.
Appendix C Additional Results on SpecBench
Methods MT-Bench Trans. Sum. QA Math RAG Avg.
[…]
Heuristics 0.82 0.83 0.85 0.83 0.83 0.83 0.83
SVIP 1.05 1.02 1.03 1.01 1.03 1.00 1.02(+56.9%)
[…]
Heuristics 1.02 0.94 0.93 0.88 1.22 0.91 0.99
SVIP 1.24 1.08 1.19 1.11 1.47 1.10 1.20(+25.0%)
[…]
Heuristics 1.56 1.55 1.76 1.49 1.61 1.55 1.58
SVIP 1.53 1.53 1.69 1.51 1.71 1.56 1.58(-1.3%)
Table 5: Speedup on SpecBench using temperature sampling.
In Figure 6, we plot the draft length and acceptance rate of Pythia (which complements Figure 5 in Section 3.1), and in Figure 7 we also give the results of the three models when using sampling instead of greedy decoding [^3] We note that in the “constant” scenario, the draft length is always set to 5. However, if an EOS token is sampled from the draft model, the draft process will terminate. So the overall draft length might be slightly lower than 5, as can be seen from Figure 7.. Across the different models and sampling methods, the same observation as discussed in Section 3 holds: SVIP results in short draft lengths and a much higher acceptance rate.
Figure 6: The average generated draft length, accepted draft length, and acceptance rate of Pythia on SpecBench.
Figure 7: The average generated draft length, accepted draft length, and acceptance rate of Qwen2.5 (top), Pythia (middle), and LLaMA-3 (bottom) on SpecBench, using temperature sampling instead of greedy decoding.
Appendix D Additional Results on Long-form Generation
Methods Generation Length
128 256 512 1K 2K 4K 6K 8K
Pythia (6.9B, ${}_{\text{(6.9B, 160M)}}$ Const. 1.10 1.30 1.50 1.66 1.82 1.20 1.11 1.04
Heuristics 1.25 1.45 1.65 1.81 2.02 1.50 1.46 1.41
SVIP 1.41 1.62 1.83 2.01 2.21 1.71 1.60 1.52
Qwen2.5 (14B, ${}_{\text{(14B, 0.5B)}}$ Const. 1.05 1.08 1.15 1.29 1.44 1.54 1.60 1.67
Heuristics 1.04 1.06 1.13 1.32 1.54 1.72 1.85 1.97
SVIP 1.30 1.34 1.42 1.57 1.74 1.87 1.98 2.10
LLaMA-3 (70B, ${}_{\text{(70B, 8B)}}$ Const. 2.06 2.18 2.31 2.45 2.58 2.72 2.77 2.78
Heuristics 2.26 2.46 2.73 3.07 3.48 3.90 4.15 4.26
SVIP 2.31 2.56 2.86 3.21 3.59 4.00 4.23 4.33
Table 6: Speedup on MT-Bench with different generation length using greedy decoding.
Figure 8: The correlation between draft model entropy and lengths of accepted draft seqeunces in the greedy setting.
In Table 6, we present the results of long-form generation with greedy decoding. Compared with Table 2, the speedup ratio of greedy decoding is much higher (even more than 4 times for LLaMA-3 after 4K context). This is due to the fact that these models tend to repeat themselves in greedy long-form generation, making it very easy for the draft models to predict the next tokens. In Figure 8, we also plot the relation between draft model entropy and accepted draft sequence lengths in this setting, which shows a much stronger correlation compared with the sampling setting in Figure 2.

### v2

@@
-Draft Model Knows When to Stop: A Self-Verification Length Policy for Speculative Decoding
+Draft Model Knows When to Stop: Self-Verification Speculative Decoding for Long-Form Generation
 Abstract
-Speculative Decoding (SD) has become an important technique in accelerating the inference speed of large language models.
-Conventional SD methods employ a fixed draft length, which ignores the token generation difficulty across tasks.
-Consequently, in this paper, we address such an issue and introduce SVIP - a difficulty-aware dynamic draft length policy for speculative decoding systems.
-Based on a theoretical lower bound of draft token acceptance rate and its inference-time approximation, SVIP adaptively determines the lengths of draft sequences based on the entropy of each draft token distribution.
-Experimental results on mainstream SD benchmarks and frameworks demonstrate the superior performance of SVIP, achieving up to 20% walltime speedup on SpecBench over baseline SD methods and 60% speedup on MT-Bench for long-form generation of up to 8K tokens.
-Moreover, SVIP is totally training-free and compatible with any existing SD methods that generate draft tokens autoregressively.
-Experimental results also show that SVIP yields consistent walltime improvement on top of GliDe & CaPE and EAGLE-2.
+Conventional speculative decoding (SD) methods utilize a predefined length policy for proposing drafts, which implies the premise that the target model smoothly accepts the proposed draft tokens.
+However, reality deviates from this assumption: the oracle draft length varies significantly, and the fixed-length policy hardly satisfies such a requirement.
+Moreover, such discrepancy is further exacerbated in scenarios involving complex reasoning and long-form generation, particularly under test-time scaling for reasoning-specialized models.
+Through both theoretical and empirical estimation, we establish that the discrepancy between the draft and target models can be approximated by the draft model’s prediction entropy: a high entropy indicates a low acceptance rate of draft tokens, and vice versa.
+Based on this insight, we propose SVIP: Self-Verification Length Policy for Long-Context Speculative Decoding, which is a training-free dynamic length policy for speculative decoding systems that adaptively determines the lengths of draft sequences by referring to the draft entropy.
+Experimental results on mainstream SD benchmarks as well as reasoning-heavy benchmarks demonstrate the superior performance of SVIP, achieving up to 17% speedup on MT-Bench at 8K context compared with fixed draft lengths, and 22% speedup for QwQ in long-form reasoning.
 Draft Model Knows When to Stop:
-A Self-Verification Length Policy for Speculative Decoding
-Ziyin Zhang1,2thanks: Work done during their internship at Tencent AI Lab.
-Jiahao Xu2 Tian Liang2 Xingyu Chen1,2∗ Zhiwei He1,2∗ Rui Wang1thanks: Corresponding author.Zhaopeng Tu2 1Shanghai Jiao Tong University 2Tencent AI Lab 1{daenerystargaryen,galaxychen,zwhe.cs,wangrui12}@sjtu.edu.cn 2{jettexu,ttianliang,zptu}@tencent.com
+Self-Verification Speculative Decoding for Long-Form Generation
+Ziyin Zhang1,2,thanks: Work done during their internship at Tencent.
+Jiahao Xu2 Tian Liang2 Xingyu Chen1,211footnotemark: 1 Zhiwei He1,211footnotemark: 1 Rui Wang1thanks: Corresponding authors.Zhaopeng Tu [^2] footnotemark: 2 1Shanghai Jiao Tong University     2Tencent 1{daenerystargaryen,galaxychen,zwhe.cs,wangrui12}@sjtu.edu.cn 2{jettexu,ttianliang,zptu}@tencent.com
+Figure 1: The variance of oracle draft length drastically increases with context length.
+MTB (MT-Bench): a conventional benchmark for SD systems.
+AIME: an extremely difficult mathematical testing set for advanced reasoning models.
 1 Introduction
 Speculative decoding (Leviathan et al., 2023; Chen et al., 2023) is a novel technique that markedly enhances the generation wall-time of large language models (LLMs).
-This approach employs a small and efficient amateur model to draft sequences, while concurrently utilizing a larger and more powerful expert model to verify the drafts.
+This approach employs a small and efficient draft model to draft sequences, while concurrently utilizing a larger and more powerful expert model to verify the drafts.
 By avoiding the autoregressive generation of each token through the target LLM, speculative decoding achieves improved efficiency while preserving the quality of the output.
-Figure 1: The “difficulty” of tokens varies in a sequence, resulting in different numbers of accepted draft tokens at different positions.
-Many variants of speculative decoding have been proposed.
-A line of work focuses on developing a stronger and faster draft model (Li et al., 2024b; Elhoushi et al., 2024; Du et al., 2024).
-Another line of work contributes to maximizing the acceptance probability of draft tokens (Sun et al., 2023; Li et al., 2024a; Lu et al., 2024).
-In general, they all tend to maximize the alignment between the draft and target model to further maximize the system acceptance rate.
-Though successful, most of these works limit their settings to a fixed draft length, where the draft model always generates a fixed number of tokens (e.g.
-4 or 5) in each iteration.
-Such a setting ignores the fact that some tokens - such as stop words or civilities - in the generation may be easy for the draft model to predict, while others - such as knowledge-intensive or reasoning-intensive tokens - can be much harder, as shown in Figure 1.
-Figure 2: The correlation between draft model entropy and draft token acceptance probability (top) and lengths of accepted draft seqeunces (bottom).
-To address this issue, in this work we introduce SVIP - Self-VerIfication length Policy, a simple, plug-and-play dynamic draft length policy for speculative decoding systems, which enhances the wall time speedup of these systems by adaptively allowing for longer draft sequences for “simple” tokens (top of Figure 1) and terminating the drafting process early upon encountering “hard” tokens (bottom of Figure 1).
-Specifically, we first analyze the acceptance rate of speculative decoding systems and derive a lower bound based on the system’s entropy information.
-Further empirical analysis suggests that such a bound can be approximated by the entropy of the draft model only, which is naturally available in the drafting process of any auto-regressive draft models.
-Consequently, we develop SVIP  which controls the length of draft sequences dynamically by determining whether to continue drafting or start verification upon the generation of each draft token.
-With extensive experiments across multiple model sizes and generation lengths, we demonstrate the superior performance of SVIP.
-It yields more than 20% of improvements over vanilla speculative decoding for Qwen2.5 14B and LLaMA-3 70B on SpecBench (Xia et al., 2024), and more than 60% of improvements for Pythia 6.9B on MT-Bench (Zheng et al., 2023) when generating long-form responses of up to 8K tokens.
-Moreover, since our method is lightweight and training-free, it is extremely flexible and can be adapted to any speculative decoding system with an auto-regressive draft model.
-As examples, we apply SVIP on top of two state-of-the-art speculative decoding methods: GliDe with a CaPE (Du et al., 2024) and EAGLE-2 (Li et al., 2024a), and confirm that it brings consistent improvements.
-Our code is available at https://github.com/Geralt-Targaryen/SVIP.
+This technique is particularly beneficial in the context of inference-time scaling, where LLMs generally generate long-form text.
+The majority of research on speculative decoding focuses on improving the acceptance rate of the draft sequences (Sun et al., 2023; Li et al., 2024b; Elhoushi et al., 2024; Du et al., 2024; Li et al., 2024a; Lu et al., 2024).
+However, they limit their settings to a fixed draft length (e.g.
+less than 5 tokens), which we find is sub-optimal in the scenario of long-form text generation.
+Figure 2: Overview of SVIP: the draft model proceeds the generation process (marked by green) until it encounters a token for which it has low confidence (marked by red), signaled by high entropy, at which point the draft model would cease generation and send the draft spans to the target model for verification.
+Specifically, as discussed in Section 2.1, we investigate the token rejection phenomenon in speculative decoding (SD) systems.
+Our findings reveal that the oracle draft length varies considerably for long-form generation, as shown in fig.˜1.
+Furthermore, heuristic methods, such as predicting the draft length, are impractically challenging because our further investigation reveals that token rejection occurs unexpectedly.
+Moreover, our closer investigation indicates a strong correlation between token rejection and the draft model’s prediction entropy at that moment.
+Inspired by such a correlation, we analyze the acceptance rate and propose our SVIP: Self-Verification Length Policy in Section 2.2.
+Specifically, we derive a lower bound for the acceptance rate based on the entropy information from the draft model.
+Notably, SVIP not only approximates this lower bound but also dynamically adjusts the length of draft sequences by determining whether to continue drafting or initiate verification after each token generation, as shown in Figure 2.
+By optimizing draft sequence lengths, SVIP enhances SD systems’ efficiency.
+Importantly, our method is entirely training-free and thus can be seamlessly integrated with any SD decoding algorithm, making it broadly applicable and efficient.
+With extensive experiments across multiple model sizes and evaluation benchmarks, we demonstrate the superior performance of SVIP  in long-context generation.
+Compared with fixed-length draft policies, it yields up to 17% speedup on MT-Bench (Zheng et al., 2023) and 22% on AIME.
+As a training-free length policy, SVIP is also extremely flexible and compatible with state-of-the-art speculative decoding systems such as EAGLE-2 (Li et al., 2024a), achieving an additional 13% speed improvement.
 In summary, our contributions are threefold:
   1. 1.
-    We derive a low bound of speculative decoding systems, where the acceptance rate of the draft model could be modeled by its entropy only.
+    We provide an in-depth analysis of the disagreement between draft model and target model in speculative decoding systems, highlighting the underperformance of fixed-length draft length policies.
   2. 2.
-    Based on this lower bound, we further develop an entropy-based dynamic draft length policy for speculative decoding systems, which is extremely flexible and can be adapted to any auto-regressive draft model.
+    Based on this analysis, we derive a low bound of speculative decoding systems, where the acceptance rate of the draft model could be modeled by its entropy only.
+We further develop SVIP, an entropy-based dynamic draft length policy for speculative decoding systems, which is extremely flexible and can be adapted to any auto-regressive draft model.
   3. 3.
-    Experimental results demonstrate the superior performance of SVIP over baseline methods, with up to 20% average speedup on SpecBench, 60% in long-form generation, and consistent improvement over state-of-the-art speculative decoding frameworks such as GliDe & CaPE and EAGLE-2.
-2 Method
-Figure 3: Distribution histograms of the entropy ratio $H_{q,p}/H_{q}$.
-For most tokens, this ratio falls into a narrow range, indicating that the cross entropy $H_{q,p}$ can be approximated by a constant multiplication of the draft entropy $H_{q}$.
-The overall objective of SVIP is to dynamically adapt draft length on-the-fly, stopping early if the current draft token’s acceptance probability is low and otherwise continuing drafting.
-To introduce our method, we first provide the background on speculative decoding in Section 2.1.
-Then, based on the key observation that the acceptance probability of a draft token depends on the target model confidence - which is unavailable in the drafting phase, we derive a theoretical lower bound for the acceptance rate based on the draft model’s entropy and the cross-entropy between the target and draft models in Section 2.2.
-Finally, based on empirical analysis of the target and draft distributions of more than 100K tokens across three model families, we approximate this lower bound using only the entropy information of the draft model in Section 2.3, making it viable in actual inference.
-2.1 Preliminaries on Speculative Decoding
-Suppose we have two LLMs $p$ and $q$, where $p$ is the larger (target) model, and $q$ is the smaller (draft) model.
-Given an input sequence $x_{\leqslant t}$ of length $t$, and a draft length $\gamma$, the draft model first samples $\gamma$ tokens $x_{t+1},\cdots,x_{t+\gamma}$ autoregressive, which are verified by the target models in parallel to acquire the confidences $p(x_{t+1}),\cdots,p(x_{t+\gamma})$ [^1] The short hand $p(x_{n})$ is used to denote the conditional probability $p(x_{n}|x_{<n})$ when there is no ambiguity.
-Throughout the work we use subscripts to indicate token indices in a sequence (e.g.
-$x_{n}$ for $n$-th token), and superscripts to indicate element indices in a vector (e.g.
-$x_{n}^{i}$ for $i$-th element in $x_{n}$)..
-[…]
-The complete algorithms for speculative decoding are given in Appendix A.
-Figure 4: Comparison between the actual acceptance probability from Equation (2), the acceptance probability lower bound from Equation (5), and the estimated lower bound after approximating the cross entropy $H_{q,p}$ with a constant multiplication of $H_{q}$.
-Each position on the x-axis corresponds to a token, which has been sorted according to the actual acceptance probability.
-2.2 Theoretical Lower Bound of Acceptance Rate
-From Section 2.1, it’s easy to derive that given an input sequence $x_{<t}$ and a draft token $x_{t}$, its acceptance probability is $\min\left(1,\;\frac{p(x_{t})}{q(x_{t})}\right)$.
-Let $\beta$ denote the expected acceptance probability over the distribution of $x_{t}$, and it follows that
-$\beta$ $=\sum_{x}q(x)\cdot\min\left(1,\frac{p(x)}{q(x)}\right)$ (1)
-$=\sum_{x}\min\left(p(x),q(x)\right).$ (2)
-Chen et al. (2023) has proven that $\beta$ is related to the total variational distance (TVD) between $p$ and $q$:
-$\beta=1-\mathrm{TVD}(p,q).$ (3)
-According to Pinsker’s inequality - which relates TVD to Kullback-Leibler divergence - we then have
-$\beta$ $\geqslant 1-\sqrt{\frac{1}{2}\mathrm{KL}(q||p)}$ (4)
-$=1-\sqrt{\frac{1}{2}\sum_{x}q(x)\log\frac{q(x)}{p(x)}}$
-$=1-\sqrt{\frac{1}{2}H_{q,p}-\frac{1}{2}H_{q}},$ (5)
+    Experimental results demonstrate the superior performance of SVIP over baseline draft length policies on both conventional long-form generation and reasoning-heavy benchmarks.
+2 Draft Model Knows When to Stop
+In this section, we first examine the behavior of draft models at the rejection phenomenon, and analyze the oracle lengths for SD systems.
+Then, we theoretically derive SVIP, which approximates the draft token acceptance rate using the draft model’s own prediction entropy.
+2.1 Investigation of Rejection
+Speculative decoding enhances the efficiency of large language model (LLM) inference by assuming that draft tokens are accepted by the LLM, thus avoiding autoregressive generation.
+Should the target model exhibit a tendency to reject tokens, the overall performance of the system may experience considerable degradation.
+Consequently, empirically investigating the rejection phenomenon is our primary interest.
+Specifically, we analyze the distribution characteristics of rejected tokens across two scenarios:
+  • AIME (2022-2024): a challenging math reasoning dataset, using greedy decoding with QwQ-32B-Preview (Team, 2024) and a 1.5B draft model;
+  • MT-Bench: a conversational and instruction-following benchmark, using sampling decoding with the Qwen2.5 family (Yang et al., 2024).
+(a) KL Div. ($\downarrow$)
+(b) Sorted vocab. distribution
+Figure 3: Agreement scores and sorted vocabulary log probability at the rejection phenomenon. The x-axis of the Figure˜3(a) represents rejected tokens and the four tokens before them.
+Rejection occurs out of the blue
+A natural question is: How does the token rejection occur? Is there any symptom before it happens? We investigate the rejected tokens with the following metrics:
+  • KL divergence for vocabulary distribution difference.
+  • Vocabulary distribution of accepted and rejected tokens.
+We quantify the occurrence of the token rejection phenomenon along with its corresponding prefix tokens, as illustrated in Figures˜3(a) and 3(b).
+It shows that KL metrics experienced a sudden and substantial surge in the position of the rejected token, and the output vocabulary distribution at rejected tokens differs significantly from previously correctly drafted tokens.
+Such a phenomenon is expected: KL divergence indicates a system discrepancy with rejection, where the ground-truth tokens are inherently difficult for the draft model to predict.
+We also quantify the entropy of the draft model at the acceptance and rejection positions in Table 1.
+It shows that the draft models suffer from a severely high entropy at rejection, indicating extreme difficulty in modeling the corresponding tokens.
+Dataset Vocab Entropy
+Accepted. Rejected
+AIME 0.25 1.30
+MT-Bench 2.18 3.99
+Table 1: Vocabulary entropy of the draft model at accepted and rejected draft tokens.
+2.2 Our Method: SVIP
+Since the rejection phenomenon occurs suddenly and the draft model suffers from high entropy at the rejection position, can we detect the rejection with KL divergence or draft model’s entropy?
+To achieve this, we seek a theoretical understanding of the rejection phenomenon.
+Lower Bound of Acceptance
+Since rejection denotes the sudden decrease of acceptance rate, we investigate the theoretic acceptance rate of the SD systems.
+Specifically, given a target model $p$, a draft model $q$, an input sequence $x_{<t}$, and a draft token $x_{t}$, it’s easy to derive that $x_{t}$’s acceptance probability is $\min\left(1,\;\frac{p(x_{t})}{q(x_{t})}\right)$ (see Appendix A).
+Let $\beta$ denote the expected acceptance probability over the distribution of $x_{t}$, and it follows that:
+$\beta$ $=\sum_{x}q(x)\cdot\min\left(1,\frac{p(x)}{q(x)}\right)$
+$=\sum_{x}\min\left(p(x),q(x)\right),$ (1)
+where $p$ and $q$ denote the target and draft model respectively.
+Chen et al.
+(2023) has proven that $\beta$ is related to the total variational distance (TVD) between $p$ and $q$.
+Start from this, we utilize Pinsker’s inequality in Equation˜3 and yield the following bound in Equation˜4:
+$\beta$ $=1-\mathrm{TVD}(p,q)$ (2)
+$\geqslant 1-\sqrt{\frac{1}{2}\mathbb{KL}(q||p)}$ (3)
+$=1-\sqrt{\frac{1}{2}H_{q,p}-\frac{1}{2}H_{q}}$ (4)
 where $H_{q,p}$ is the cross entropy between $q$ and $p$, and $H_{q}$ is the entropy of $q$.
-Equation (5) provides a theoretical lower bound for the acceptance rate of a draft token using the entropy information of the speculative decoding system [^2] An alternative way to construct the lower bound is discussed in Appendix B..
-To provide an intuitive motivation for using entropy to construct the lower bound, we plot the relation between draft model entropy and draft token acceptance probability as well as accepted draft seqeunce lengths in Figure 2, which are collected from more than 1M tokens generated by three different target models with temperature set to 1 and maximum draft length set to 40.
-The first subfigure indicates a strong negative correlation between draft model entropy and draft token acceptance probability, while the second subfigure shows that when the entropy is low, dozens of consecutive draft tokens could be accepted, which highlights the drawback of setting draft length to a constant, small value, as is the common practice in speculative decoding literature.
-Another version of Figure 2 in the greedy setting is given in Appendix D.
-2.3 Empirical Estimation of Acceptance Rate
-So far in this section, we have been assuming access to the target distribution $p(x_{t})$ of the next token, which is unavailable in the drafting phase at inference time.
-Thus, to apply Equation (5) for actual acceptance rate estimation, we must approximate the cross entropy $H_{q,p}$ with information from only the draft model.
+We denote the above bound as the oracle bound.
+Utilizing this bound for acceptance prediction is infeasible since it requires instantaneous access to the target model for cross entropy $H_{q,p}$, which is infeasible during the drafting phase.
+Approximating the Oracle Bound with Draft Distribution
+Can we approximate the cross-entropy $H_{q,p}$ between the draft model’s distribution $q$ and the target distribution $p$ using only $q$?
+We propose using the draft model’s entropy $H_{q}$ as a proxy, approximating $H_{q,p}$ as $\gamma H_{q}$, where $\gamma=H_{q,p}/H_{q}$ is a random variable capturing the ratio between $H_{q,p}$ and $H_{q}$.
+This leads to a bound on the acceptance rate $\beta$:
+$\beta\geq 1-\sqrt{\frac{1}{2}(\gamma-1)H_{q}}.$ (5)
+To make this bound practical, we approximate $\gamma$ with a constant $c$, yielding the approximation bound:
+$\beta\geq 1-\sqrt{\frac{1}{2}(\gamma-1)H_{q}}\approx 1-\sqrt{cH_{q}}.$ (6)
+This bound holds when $1-\sqrt{\frac{1}{2}(\gamma-1)H_{q}}\geq 1-\sqrt{cH_{q}}$, ensuring the approximation is conservative.
+Thus, the approximation bound $1-\sqrt{cH_{q}}$ lower-bounds the true acceptance rate $\beta$.
+Detecting Rejection with Draft Entropy
+With Equation˜6 providing a way to estimate the acceptance probability using only the draft model’s entropy, we introduce SVIP, which dynamically adapts the draft length.
+After generating each draft token, we compute the approximation bound and halt drafting if it falls below a threshold $\hat{h}$, i.e., if $1-\sqrt{cH_{q}}<\hat{h}$.
+Since $c$ and $\hat{h}$ are constant hyperparameters, we simplify the criterion by defining a new threshold $h=(1-\hat{h})/\sqrt{c}$, absorbing $\sqrt{c}$ into $h$.
+This reduces the stopping condition to $\sqrt{H_{q}}>h$.
+Formally, given a prefix $x_{<t}$ of t1t-1 tokens, the stopping criterion is:
+$\sqrt{H_{q}(x_{<t})}>h,$ (7)
+where $H_{q}(x_{<t})$ is the entropy of the draft distribution conditioned on $x_{<t}$.
+This ensures drafting stops when the estimated acceptance probability is too low, optimizing efficiency while maintaining reliability.
+We formalize SVIP in Algorithm 1.
+The details of the methods Verify and Correct are given in Appendix A, for which different versions are available for sampling (Algorithm 2, 4) and greedy decoding (Algorithm 3, 5).
 Algorithm 1 SVIP
@@
 15:  for $j=1$ to $\gamma$ do
-16:         if Verify$\left(p_{x|x_{<n+j}},\,q_{x|x_{<n+j}},x_{n+j}\right)$ then
+[…]
 17:     $\tilde{n}\leftarrow\tilde{n}+1$
@@
 27:end while
-28:$x_{\leqslant n}$
-To tackle this issue, we first plot the relationship between $H_{q,p}$ and $H_{q}$ in Figure 3.
-For all three model families, the entropy ratio $H_{q,p}/H_{q}$ are concentrated in a narrow range between 1.0 and 1.3.
-Thus, we choose to approximate $H_{q,p}$ with a constant multiplication of $H_{q}$.
-Plugging it into Equation (5), we now have
-$\beta$ $\geqslant 1-\sqrt{cH_{q}},$ (6)
-where $c$ is a constant controlling the approximation ratio between $H_{q,p}$ and $H_{q}$. In Figure 4, we visualize the values derived from Equation (2), (5), and (6).
-With Equation (6) providing a way to estimate acceptance probability using only the draft model’s entropy, we can now adapt the draft length on-the-fly.
-After generating each draft token, we compute the estimated acceptance probability lower bound, and stop the draft process if it’s lower than a certain threshold $h$.
-We note that since both $c$ and $h$ are constant hyperparameters, we can remove $\sqrt{c}$ from Equation 6 and absorb it into the threshold $h$.
-We formalize SVIP in Algorithm 1.
-The details of the methods Verify and Correct are given in Appendix A, for which different versions are available for sampling (Algorithm 2, 4) and greedy decoding (Algorithm 3, 5).
+[…]
+Figure 4: Comparison between the actual acceptance probability in Equation˜1, the acceptance probability lower bound in Equation˜4, and the estimated lower bound in Equation˜6.
+Each position on the x-axis corresponds to a token, which has been sorted according to the actual acceptance probability.
+Figure 5: The SD system speedup on MT-Bench using Qwen2.5-14B (top) and Qwen2.5-32B (bottom) as targets and three different smaller models as drafts.
+Justifying the Approximation Bound
+The tightness of the approximation depends on how well $c$ captures the behavior of the random variable $\gamma$. For the approximation bound to be conservative, it requires:
+$\gamma\leq 2c+1.$ (8)
+Given the right-skewed nature of $\gamma\geq 1$, we model $\gamma=1+X$, where $X\sim\text{Gamma}(\alpha,\beta)$. The probability that the approximation is valid is:
+$P(\gamma\leq 2c+1)=P(X\leq 2c)=\frac{\gamma(\alpha,\beta\cdot 2c)}{\Gamma(\alpha)},$ (9)
+where $\gamma(\alpha,z)=\int_{0}^{z}t^{\alpha-1}e^{-t}\,dt$ is the lower incomplete gamma function, and $\Gamma(\alpha)$ is the gamma function. The choice of $c$ trades off reliability and tightness:
+  • If $c$ is small (e.g., $2c<\mathbb{E}[X]=\alpha/\beta$), the probability $P(X\leq 2c)$ is low, risking an invalid bound.
+  • If $c$ is large, $P(X\leq 2c)\to 1$, but the bound $\beta\geq 1-\sqrt{cH_{q}}$ becomes looser.
+Optimal performance requires balancing the reliability of the approximation (high $P(X\leq 2c)$) with the tightness of the bound (small $c$).
+We analyze such a trade-off on Qwen2.5 on MT-Bench and QwQ-32B on AIME in Figure˜4.
+It shows that for most cases our estimated approximation bound works well (has a higher acceptance probability than the oracle bound when hyperparameter $c$ (i.e.
+$h$) is properly selected, set to 0.18 in the figure).
 3 Experiments
-3.1 Experiments on SpecBench
-3.1.1 Settings
-Methods MT-Bench Trans. Sum. QA Math RAG Avg.
-Pythia (6.9B, ${}_{\text{(6.9B, 160M)}}$ Const. 1.45 1.47 1.24 1.43 1.52 1.42 1.42
-Heuristics 1.51 1.58 1.34 1.58 1.64 1.51 1.53
-SVIP 1.63 1.62 1.45 1.67 1.72 1.66 1.63(+14.8%)
-Qwen2.5 (14B, ${}_{\text{(14B, 0.5B)}}$ Const. 1.08 0.87 1.11 0.92 1.43 0.99 1.07
-Heuristics 1.10 0.91 1.10 0.92 1.34 1.03 1.07
-SVIP 1.33 1.12 1.37 1.14 1.57 1.23 1.29(+20.6%)
-LLaMA-3 (70B, ${}_{\text{(70B, 8B)}}$ Const. 2.04 2.48 2.56 2.34 2.32 2.28 2.34
-Heuristics 2.30 3.13 3.33 2.61 2.52 2.63 2.76
-SVIP 2.31 3.04 3.48 2.63 2.89 2.59 2.83(+20.9%)
-Table 1: Speedup over target-model-only autoregressive decoding on SpecBench.
-We validate the effectiveness of SVIP on Spec-Bench (Xia et al., 2024) using three distinct target models: Pythia 6.9B (Biderman et al., 2023), Qwen2.5 14B (Yang et al., 2024), and LLaMA-3 70B (Dubey et al., 2024), with Pythia-160M, Qwen2.5 0.5B, and LLaMA-3 8B as the draft models respectively.
-As baselines, we consider two simple policies for draft length: 1) a constant draft length of 5, which is commonly used in the literature, and 2) the heuristics implemented in Hugging Face Transformers library (Wolf et al., 2019), where the draft length for the next draft iteration is increased by 2 if all draft tokens in the current iteration are accepted, and otherwise decreased by 1.
-We set the sampling temperature to 0 on SpecBench (the alternatives are discussed in Appendix C).
-For each model, the entropy threshold $t$ in SVIP is chosen from $\{0.2,0.3,0.4,0.5\}$ based on performance on 8 samples held out from MT-Bench (Zheng et al., 2023), which are 0.4 for Pythia, and 0.3 for Qwen2.5 and LLaMA-3.
-All experiments with Pythia and Qwen are conducted on a single 40GB A100, while experiments with LLaMA are conducted on 5 40GB A100s.
-To mitigate the impact of system performance varitations, we repeat all experiments with Pythia and Qwen for three times (using different random seeds when they are used) and report the average speedup over target-model-only autoregressive decoding.
+Next, to verify the effectiveness of SVIP, we conduct experiments on both conventional long-form generation (Section 3.1) and reasoning with test-time scaling (Section 3.2).
+Since SVIP is completely training-free, we also apply it to other speculative decoding methods and demonstrate its flexibility (Section 3.1).
+As baselines, we consider two widely adopted policies for draft length:
+  1. 1.
+    Constant: a constant draft length (set to 5 unless otherwise stated), which is commonly used in the literature
+  2. 2.
+    Heuristic: the heuristics implemented in Hugging Face Transformers library (Wolf et al., 2019), where the draft length for the next draft iteration is increased by 2 if all draft tokens in the current iteration are accepted, and otherwise decreased by 1.
+3.1 Results on Long-form Generation
+Settings
+We first validate the effectiveness of SVIP on the widely used MT-Bench (Zheng et al., 2023) using different sizes of Qwen2.5 (Yang et al., 2024) as target and draft models.
+Unlike many existing works on speculative decoding (Chen et al., 2023; Du et al., 2024) that limit their experiments to generating short sequences of 128 tokens, we conduct experiments on long-form generation with up to 8K context to investigate the applicability of speculative decoding in a broader scope.
+We set the sampling temperature to 1, as we found that when using greedy decoding in long-form generation, both the draft and the target models are prone to repeat themselves, resulting in very low information entropy and exaggerated speedup ratios (Ouyang et al., 2024) (see Appendix D for details).
+The entropy threshold $h$ in SVIP is chosen from $\{0.2,0.3,0.4,0.5\}$ based on performance on 8 held out samples using the 14B model as target and the 0.5B model as draft, which is set to 0.3 and reused in all following experiments.
+As evaluation metrics, we mainly report the average speedup over target-model-only autoregressive decoding, but also consider other auxiliary information including accepted draft lengths and draft token accept rate.
 Also, since the memory consumption of verifying $n$ draft tokens is quadratic in $n$, we limit the maximum draft length to 40 in both heuristics and SVIP scenarios, beyond which we start to encounter out-of-memory issues.
-3.1.2 Results
-The results on SpecBench are shown in Table 1.
-Compared with the constant approach, SVIP yields an average speedup of 15% for Pythia and 20% for Qwen and LLaMA.
-Compared with the heuristic approach, SVIP also gives a consistent improvement on all domains for Pythia and Qwen, and outperforms the latter on 4 out of 6 domains for LLaMA.
-Figure 5: The average generated draft length, accepted draft length, and acceptance rate of Qwen2.5 (top) and LLaMA-3 (bottom) on SpecBench.
-Compared with the two baselines, SVIP leads to a shorter draft length and a much higher acceptance rate.
-In Figure 5, we plot the average draft length and accepted draft length of Qwen and LLaMA (the results for Pythia are similar, and are given in Appendix C).
-From the figure, we observe that by terminating the draft process when the draft model entropy is high, SVIP leads to shorter draft lengths and a much higher acceptance rate (close to or more than 80% on all domains for Qwen, and more than 90% for LLaMA).
-Notably, the acceptance rate of LLaMA-3 even reaches 99% on the summarization domain, which contributes to the highest speedup (3.48) in Table 1.
-3.2 Long-form Generation
-Most existing works on speculative decoding (Chen et al., 2023; Du et al., 2024) limit their experiments to generating short sequences of 128 tokens.
-To verify the wide applicability of SVIP, we also conduct experiments on long-form generation with up to 8K context.
-For this purpose, we use MT-Bench (Zheng et al., 2023) as the dataset and set the sampling temperature to 1, as we found that when using greedy decoding in long-form generation, both the draft and the target models are prone to repeat themselves, resulting in very low information entropy (see Appendix D for details).
-Other settings follow Section 3.1.
-Methods Generation Length
-128 256 512 1K 2K 4K 6K 8K
-Pythia (6.9B, ${}_{\text{(6.9B, 160M)}}$ Const. 0.68 0.69 0.69 0.70 0.72 0.90 0.89 0.87
-Heuristics 0.88 0.88 0.88 0.88 0.90 1.21 1.25 1.23
-SVIP 1.07 1.08 1.08 1.07 1.08 1.43 1.44 1.41
-Qwen2.5 (14B, ${}_{\text{(14B, 0.5B)}}$ Const. 0.98 0.97 0.95 0.96 0.98 1.00 1.02 1.04
-Heuristics 1.01 0.99 0.98 1.00 1.02 1.03 1.04 1.06
-SVIP 1.29 1.29 1.30 1.31 1.32 1.33 1.33 1.35
-LLaMA-3 (70B, ${}_{\text{(70B, 8B)}}$ Const. 1.74 1.74 1.77 1.78 1.82 1.90 1.93 1.94
-Heuristics 1.53 1.56 1.61 1.63 1.68 1.77 1.83 1.84
-SVIP 1.69 1.72 1.75 1.78 1.86 1.96 2.01 2.02
-Table 2: Speedup on MT-Bench with different generation length.
-The results are given in Table 2.
-Interestingly, we find that in the sampling setting, the two baseline methods (constant and heuristics) perform even worse than target-model-only auto-regressive decoding for Pythia and Qwen, while SVIP consistently yields a positive speedup.
-For LLaMA, while constant draft length performs better with contexts shorter than 1K tokens, SVIP exceeds it for generating longer sequences.
-Another observation is that the speedup ratio of all three methods generally increases with the context length, which could be possibly attributed to the longer contexts giving the draft model more information, aligning it at test time to the target model.
-However, we note that the absolute values of token throughput (which are not discussed in this paper, since they heavily depend on the underlying machines) stay at the same level from 1K to 8K context for Qwen2.5 and LLaMA-3 when using speculative decoding, and decrease slowly when not using speculative decoding.
-For Pythia, which does not use any optimized attention such as MQA (Shazeer, 2019) or GQA (Shazeer, 2019), the throughput decreases notably with context length in both cases - with or without specualtive decoding.
-3.3 Applying SVIP to Other Draft Methods
-In Section 3.1 and 3.2, we evaluated SVIP on vanilla speculative decoding, where a standard pretrained Transformer decoder model from the target model’s family is used as the draft model.
+SVIP outperforms all baselines
+We validate our proposed method, SVIP, with two target models: Qwen2.5-14B and Qwen2.5-32B, utilizing draft models that vary in size from 0.5B to 3B.
+We show the performance of SVIP and baselines in Figure 5.
+It shows that SVIP consistently outperforms constant and heuristics draft length by a large margin.
+Figure 6: Analysis of Qwen2.5 14B/0.5B’s behaviours on MT-Bench. Compared with constant and heuristics length policies, SVIP generates shorter drafts with a significantly higher accept rate.
+Figure 7: Further analysis of Qwen2.5 14B/0.5B on MT-Bench: (a) SVIP outperforms all constant length policies ranging from 2-5; (b) delta draft length (defined as proposed draft length minus oracle draft length) show that constant and heuristics length policies tend to over-generate drafts, while SVIP models the oracle draft length almost perfectly.
+To further investigate the origin of SVIP’s speedup, we analyze the proposed draft lengths and accepte rate of the different length policies in Figure 6.
+We observe that by terminating the draft process early when the draft model entropy is high, SVIP leads to shorter draft lengths and a much higher acceptance rate.
+However, in Figure 7 we also compare the performance of SVIP with shorter constant draft length policies, which suggest that simply using shorter constant draft length does not suffice, highlighting the importance of dynamically determining draft lengths.
+Figure 6 also suggests that while the heuristics draft length policy tends to produce very long drafts at long context, its acceptance rate remains low, resulting in no effective speedup compared with fixed draft length.
+SVIP  drafts fit oracle length
+In Figure 7, we revisit the concept of “oracle draft length” introduced in Section 2.1, and plot the differences between actual draft lengths and oracle draft lengths.
+The results suggest that both constant and heuristics draft length policies tend to generate drafts that are too long, while SVIP models the oracle draft length almost perfectly, with an average delta below 0.5 tokens.
+SVIP further boosts strong baseline
+In the previous experiments, we evaluated SVIP on vanilla speculative decoding, where a standard pretrained Transformer decoder model with the same vocabulary as the target is used as the draft model.
 However, in the past years many works on speculative decoding have proposed other stronger or more efficient draft models (Cai et al., 2024; Du et al., 2024; Li et al., 2024b).
 Since most of these works assume a constant draft length, SVIP is orthogonal to them and can be applied on top of them without any additional training.
-Methods MT-Bench Code Finance GSM Spider Avg.
-7B GliDe 1.95 2.04 1.91 1.98 1.69 1.95
-   +SVIP 2.00 2.12 2.03 2.01 1.63 2.02
-GliDe + CaPE 2.36 2.57 2.29 2.51 1.97 2.40
-   +SVIP 2.56 2.65 2.49 2.54 2.08 2.52
-13B GliDe 2.22 2.41 2.15 2.31 1.85 2.24
-   +SVIP 2.31 2.43 2.17 2.35 1.85 2.28
-GliDe + CaPE 2.73 2.86 2.66 2.80 2.24 2.73
-   +SVIP 2.72 2.93 2.66 2.85 2.27 2.76
-33B GliDe 2.12 2.25 2.09 2.29 1.99 2.18
-   +SVIP 2.29 2.40 2.20 2.42 2.03 2.30
-GliDe + CaPE 2.08 1.98 2.10 2.13 1.76 2.03
-   +SVIP 2.13 2.02 2.15 2.16 1.82 2.08
-Table 3: Speedup comparison with GliDe & CaPE, using Vicuna as the base model.
-Specifically, we consider GliDe with a CaPE (Du et al., 2024), where the draft model - named GliDe - is a small transformer decoder with cross-attention to the target model’s hidden representations, while CaPE is a complementary tree expansion method to increase the acceptance rate of the draft token at each position.
-Following the settings of Du et al.
-(2024), we use Vicuna 7B, 13B, and 33B (Chiang et al., 2023) as the base models (for which the draft models are publicly available) and set the sampling temperature to 0.
-However, we distribute the 33B model across two GPUs due to memory constraint.
-Similar to the previous experiments, we set the threshold $t$ to 0.5 based on pilot experiments on 8 samples held out from MT-Bench using GliDe only (without CaPE).
-We also apply SVIP to EAGLE-2 (Li et al., 2024a), the state-of-the-art speculative decoding system which utilizes the target model’s language modeling head on top of the draft model’s features to predict the next draft token, and dynamically constructs a draft tree at each draft position.
+Table 2: Speedup on MT-Bench on top of EAGLE-2, using Vicuna as base models.
+Model Method Context Length
+128 256 512 1K 2K 4K
+V-7B E2 2.70 2.62 2.52 2.41 2.43 1.24
++ SVIP 2.80 2.76 2.71 2.69 2.75 1.41
+V-13B E2 2.95 2.90 2.83 2.74 2.71 1.53
++ SVIP 2.94 2.99 2.93 2.86 2.79 1.64
+Specifically, we also apply SVIP to EAGLE-2 (Li et al., 2024a), the current state-of-the-art (SOTA) speculative decoding system which utilizes the target model’s language modeling head on top of the draft model’s features to predict the next draft token, and dynamically constructs a draft tree at each draft position.
 Following Li et al.
-(2024a), we use LLaMA-2 7B, 13B (Touvron et al., 2023) and Vicuna 7B, 13B (Chiang et al., 2023) as the base models, and set the sampling temperature to 1.
-As Brown et al.
-(2024) suggest, adding a conditional clause (that decides whether or not to stop drafting) after the generation of each draft token in EAGLE-2 may interrupt the otherwise input-independent control flow of EALGE-2, reducing the effects of low-level interpreter and system optimizations.
-Thus, in EAGLE-2 we calculate the draft entropy and decide whether or not to stop drafting after every two draft tokens instead of after every draft token, as used in previous experiments.
-Methods MT-Bench H-Eval GSM8K Alpaca CNN/DM QA Avg.
-LLaMA-2 7B EAGLE-2 3.10 3.61 3.15 3.10 2.76 2.84 3.10
-   + SVIP 3.16 3.66 3.18 3.13 2.82 3.02 3.16
-LLaMA-2 13B EAGLE-2 3.38 4.12 3.41 3.25 3.01 2.98 3.41
-   + SVIP 3.41 4.09 3.45 3.34 3.05 3.17 3.46
-Vicuna 7B EAGLE-2 2.66 2.84 2.77 2.48 2.31 2.13 2.58
-   + SVIP 2.84 2.97 2.75 2.66 2.42 2.30 2.73
-Vicuna 13B EAGLE-2 2.85 3.31 2.93 2.74 2.48 2.30 2.83
-   + SVIP 2.94 3.49 3.19 2.89 2.60 2.53 2.98
-Table 4: Speedup comparison with EAGLE-2, using LLaMA-2-Chat and Vicuna as the base models.
-The results for these two methods are presented in Table 3 and 4, respectively.
-For GliDe, SVIP yields a consistent speedup both with and without CaPE, with a 5% improvement for the 7B and 33B models.
-For EAGLE-2, the speedup is also particularly notable for the Vicuna models, with 5% improvement for both the 7B and 13B models.
+(2024a), we use Vicuna 7B, 13B (Chiang et al., 2023) as the base models, and set the sampling temperature to 1.
+To the best of our knowledge, we are also the first to investigate EAGLE-2’s effectiveness in long-form generation.
+The results of EAGLE-2 on MT-Bench are given in Table 2.
+Even on top of this SOTA speculative decoding system, SVIP yields consistent improvement, which is especially notable at longer context length, surpassing the vanilla EAGLE-2 by 14% speedup for Vicuna 7B and 7% for Vicuna 13B.
+Table 3: Speedup of QwQ on MATH, AIME, and GPQA, along with their average generation length.
+MATH500 GPQA AIME Avg
+Level1 Level2 Level3 Level4 Level5
+Avg. Length 1.3K 1.3K 1.6K 2.5K 3.6K 3.9K 6.2K
+Const. 1.45 1.50 1.52 1.56 1.56 1.25 1.58 1.49
+Heuristics 1.29 1.26 1.27 1.30 1.33 1.18 1.34 1.28
+SVIP 1.65 1.68 1.75 1.78 1.82 1.52 1.77 1.71
+3.2 Long-form Reasoning
+Settings
+Recently, o1-style reasoning models have come into the spotlight of LLM research.
+Thus, we are especially interested in seeing the effectiveness of SVIP and other speculative decoding strategies on such models, which often have very long outputs.
+Consequently, we utilize QwQ-32B-Preview (Team, 2024), the only applicable open-source reasoning model at the time of writing, which does not have off-the-shelf smaller variants, so we train our own draft model based on Qwen2.5 1.5B by distilling QwQ 32B on 1M mathematical Persona data (Chan et al., 2024) [^1] We have made this model publicly available and will provide links in camera-ready version..
+Using this draft model, we conduct experiments on MATH (Hendrycks et al., 2021), AIME [^2] https://maa.org/maa-invitational-competitions/, and GPQA (Rein et al., 2023).
+We sample 200 questions from MATH ranging from level 1 to level 5, and use 73 questions released from 2022 to 2024 for AIME.
+For GPQA, we use the diamond test set.
+SVIP achieves strong speedup in long-form reasoning
+The overall results are given in Table 3.
+SVIP outperforms the two baselines by a large margin across different benchmarks and context lengths.
+Detailed analysis of the different length policies’ behaviours suggests similar results to the previous experiment: SVIP has an average proposal length similar to the constant draft length policy, but with a much higher draft token accept rate, leading to more effective speedup.
+Figure 8: A case comparison of different length policies during one generation example.
+Drafts proposed by SVIP vary drastically in length, while constant and heuristics policies are insufficient to model such changes.
+Table 4: The average draft model entropy and draft token acceptance rate of some representative tokens between QwQ-32B and 1.5B.
+Token Avg. Entropy Accept Rate
+All tokens 0.38 0.68
+“Wait” 1.17 0.53
+“Alright” 1.38 0.22
+“Actually” 1.52 0.33
+“)]” 0.12 1.00
+“}}” 0.02 1.00
+“ynomials” 0.01 1.00
+“ponents” 0.01 1.00
+In Table 4, we present some token cases with very low or very high acceptance rate.
+We find that completing subword units or equations is quite easy for the small draft model, while several keywords in QwQ’s reasoning patterns are much harder.
+The reverse correlation of draft entropy and acceptance rate on these tokens further validates the motivation of SVIP.
+In Figure 8, we plot the behaviors of different draft length policies in an example generation case.
+The drastic draft length oscillations in SVIP also highlight the importance of dynamic draft length policy.
 4 Related Work
@@
 Some works also take the best of both worlds and introduce extra layers on top of the target model’s hidden representations to construct draft models, represented by EAGLE (Li et al., 2024b), GliDe (Du et al., 2024), and Medusa (Cai et al., 2024).
+Unlike previous methods, SVIP  has no requirement for draft models except for autoregression, and is a totally training-free adaptive-length policy, which could boost any draft model’s performance.
 Draft tree expansion.
@@
 Other works have also addressed the problem of multi-draft verification from a theoretic perspective (Sun et al., 2023; Yin et al., 2024).
+In contrast to previous methods which introduce tree expansion for proposing fixed fine-grained n-gram draft, our method SVIP involves no fixed tree expansion rules, and is mainly about a more dynamic and flexible draft length policy.
 Draft length control.
@@
 Brown et al.
-(2024) propose Dynamic Depth Decoding on top of EAGLE-2, which uses the sum of all tokens’ confidences in one level of its draft tree as an indicator to predict whether or not to continue draft generation.
+(2024) propose Dynamic Depth Decoding (DDD) on top of EAGLE-2, which uses the sum of all tokens’ confidences in one level of its draft tree as an indicator to predict whether or not to continue draft generation.
+Concurrent with our work, Zhang et al.
+(2024) propose AdaEAGLE, which utilizes an MLP on top of EAGLE to predict the next round’s draft length.
+In contrast to prior length-control strategies that necessitate the training of a length-prediction module, SVIP stands out with its training-free nature.
+This unique characteristic endows it with remarkable flexibility, allowing it to be seamlessly integrated and applied to any autoregressive draft model.
 5 Conclusion
@@
 With extensive experiments spanning various base models, draft methods, test domains, and generation length, we validated the effectiveness of SVIP, sparking new insights on speculative decoding and more efficient large language models.
+For future work, we aim to investigate tighter bounds on the acceptance rate to improve the accuracy of acceptance probability estimates, thereby enabling more efficient draft length adaptation.
+Additionally, our current analysis may not fully capture context-dependent patterns.
+A more nuanced investigation into these patterns could further enhance performance.
+Limitations
+While SVIP advances speculative decoding through adaptive draft length control, it has several limitations that offer avenues for future work.
+The acceptance rate bound in Equation˜6 could be overly conservative.
+Developing a tighter bound would enhance the accuracy of acceptance probability estimates, enabling more effective draft length adaptation.
+Further, our analysis assumes they follow a simplified distribution discrepancy of SD systems.
+This may not fully capture the nuanced factors contributing to their occurrence, such as context-dependent patterns or model-specific biases.
+Context-dependent length proxy for SD systems could be the potential research direction.
 Appendix A The Complete Speculative Decoding Algorithms
@@
 1:target distribution $p(x)$, draft distribution $q(x)$, draft token $x_{t}$
-2:$accept\leftarrow$ False
-3:$r\sim U[0,1]$
+[…]
+[…]
 4:if $r<\frac{p(x_{t})}{q(x_{t})}$ then
@@
 6:end if
-7:$accept$
+7:acceptaccept
 Algorithm 3 Verify (Greedy)
 1:target distribution $p(x)$, draft distribution $q(x)$, draft token $x_{t}$
-2:$accept\leftarrow$ False
+[…]
 3:if $\arg\max p(x)==x_{t}$ then
@@
 5:end if
-6:$accept$
+6:acceptaccept
 Algorithm 4 Correct (Sampling)
 1:target distribution $p(x)$, draft distribution $q(x)$
+2:Sample $\hat{x}\sim\frac{\max(q(x)-p(x),\,0)}{\sum_{i}\max(q(x^{i})-p(x^{i}),\,0)}$
 […]
-3:$\hat{x}$
 Algorithm 5 Correct (Greedy)
 1:target distribution $p(x)$, draft distribution $q(x)$
-2:$\arg\max p(x)$
+[…]
 Algorithm 6 Speculative Decoding
@@
 9:  for $j=1$ to $\gamma$ do
-10:         if Verify$\left(p(x|x_{<n+j}),\,q(x|x_{<n+j}),x_{n+j}\right)$ then
+[…]
 11:     $\tilde{n}\leftarrow\tilde{n}+1$
@@
 21:end while
-22:$x_{\leqslant n}$
+[…]
+We note that from Algorithm 2, it’s straightforward that the acceptance rate of a draft token $x_{t}$ in the sampling scenario is by definition $\min(1,\frac{p(x_{t})}{q(x_{t})})$.
 Appendix B Alternatives for Acceptance Rate Lower Bound Computation
@@
 $\beta$ $=\sum_{x}\min\left(p(x),q(x)\right)$ (10)
-$\geqslant 1-\sqrt{\frac{1}{2}KL(q||p)}.$ (8)
+$\geqslant 1-\sqrt{\frac{1}{2}\mathbb{KL}(q||p)}.$ (11)
 Another way to compute the lower bound of acceptance probability can be derived from Bretagnolle-Huber inequality (Bretagnolle and Huber, 1978):
-$\beta$ $\geqslant 1-\sqrt{1-e^{-KL(q||p)}}.$ (9)
-Compared with the Pinsker’s bound, it’s trivial to see that this bound is guaranteed to be always larger than 0.
-However, in practice we find that the Pinsker’s bound is 11% tighter for Qwen2.5, 20% tighter for Pythia, and 43% tighter for LLaMA-3.
-Appendix C Additional Results on SpecBench
-Methods MT-Bench Trans. Sum. QA Math RAG Avg.
-[…]
-Heuristics 0.82 0.83 0.85 0.83 0.83 0.83 0.83
-SVIP 1.05 1.02 1.03 1.01 1.03 1.00 1.02(+56.9%)
-[…]
-Heuristics 1.02 0.94 0.93 0.88 1.22 0.91 0.99
-SVIP 1.24 1.08 1.19 1.11 1.47 1.10 1.20(+25.0%)
-[…]
-Heuristics 1.56 1.55 1.76 1.49 1.61 1.55 1.58
-SVIP 1.53 1.53 1.69 1.51 1.71 1.56 1.58(-1.3%)
-Table 5: Speedup on SpecBench using temperature sampling.
-In Figure 6, we plot the draft length and acceptance rate of Pythia (which complements Figure 5 in Section 3.1), and in Figure 7 we also give the results of the three models when using sampling instead of greedy decoding [^3] We note that in the “constant” scenario, the draft length is always set to 5.
-However, if an EOS token is sampled from the draft model, the draft process will terminate.
-So the overall draft length might be slightly lower than 5, as can be seen from Figure 7..
-Across the different models and sampling methods, the same observation as discussed in Section 3 holds: SVIP results in short draft lengths and a much higher acceptance rate.
-Figure 6: The average generated draft length, accepted draft length, and acceptance rate of Pythia on SpecBench.
-Figure 7: The average generated draft length, accepted draft length, and acceptance rate of Qwen2.5 (top), Pythia (middle), and LLaMA-3 (bottom) on SpecBench, using temperature sampling instead of greedy decoding.
+$\beta$ $\geqslant 1-\sqrt{1-e^{-\mathbb{KL}(q||p)}}.$ (12)
+Compared with the Pinsker’s bound, it’s trivial to see that this bound is guaranteed to be always larger than 0. However, in practice we find that the Pinsker’s bound is about 11% tighter.
+Appendix C $\gamma$ Approximation
+C.1 Approximation Bound
+Following Eq. (4), the acceptance rate $\beta$ satisfies:
+$\beta\geq 1-\sqrt{\frac{1}{2}\mathbb{KL}(q||p)}=1-\sqrt{\frac{1}{2}H_{q,p}-\frac{1}{2}H_{q}}$
+We denote the above bound as the actual bound.
+While this bound is theoretically sound, it relies on the exact access to the $H_{q,p}$ the cross entropy between target and the draft models, which is inaccessible during the SD drafting phase.
+To address this, approximate $H_{q,p}$ with $\gamma H_{q}$, where $\gamma$ is a random variable to describe the ratio between $H_{q,p}$ and $H_{q}$, i.e.
+$\gamma=H_{q,p}/H_{q}$, we could rewrite the bound:
+$\mathbb{KL}(q||p)=H_{q,p}-H_{q}=(\gamma-1)H_{q}$
+$\beta\geq 1-\sqrt{\frac{1}{2}(\gamma-1)H_{q}}$
+To make this bound more practical, we approximate it using a constant $c$, obtaining
+$\beta\geq 1-\sqrt{\frac{1}{2}(\gamma-1)H_{q}}\approx 1-\sqrt{cH_{q}}$
+We denote the above bound as the approximation bound.
+Since $\gamma$ is a random variable, the tightness and reliability of this approximation depend on how well $c$ aligns with $\gamma$’s behavior.
+Specifically, we need our approximation bound is smaller than the actual bound:
+$\beta\geq 1-\sqrt{\frac{1}{2}(\gamma-1)H_{q}}\geq 1-\sqrt{cH_{q}}$
+Simplify the right side inequality:
+$\gamma\leq 2c+1$ (13)
+C.2 Theoretical Analysis
+Now, from the $\gamma$’s distribution in Figure 5, let’s analyze the probability that the lower bounds hold by modeling $\gamma$’s distribution.
+Gaussian Distribution
+Let’s assume $\gamma\sim N(\mu,\sigma^{2})$, and the probability that the bound holds is:
+$P(\gamma\leq 2c+1)=\Phi(\frac{2c+1-\mu}{\sigma})$ (14)
+where $\Phi$ is the standard normal CDF. It demonstrates that:
+  • If $c$ is small (e.g. $2c+1\leq\mu$), the probability that the bound holds is low.
+  • If $c$ is large (e.g. $2c+1\geq\mu$), the bound holds with high probability; however the bound itself becomes loose.
+Gamma Distribution
+Given the right-skewed nature of $\gamma\geq 1$, we model as a shifted Gamma distribution: $\gamma=1+X$, where $X\sim\text{Gamma}(\alpha,\beta)$. The conditions for the bound to hold is :
+$P(\gamma\leq 2c+1)=P(X\leq 2c)=\frac{\gamma(\alpha,\beta\cdot 2c)}{\Gamma(\alpha)}$ (15)
+where $\gamma(\alpha,z)=\int_{0}^{z}t^{\alpha-1}e^{-t}dt$ is the lower incomplete gamma function, and $\Gamma(\alpha)$ is the gamma function. This probability depends on $c$, $\alpha$, and $\beta$.
+  • If $c$ is small (e.g. $2c<\mathbb{E}[x]=\alpha/\beta$), so the probability that our approximation further lowers the actual bound is low.
+  • It $c$ is large, the probability approaches 1, but the bound $\beta\geq 1-\sqrt{cH_{q}}$ is looser.
 Appendix D Additional Results on Long-form Generation
-Methods Generation Length
+In Table 5, we present the results of greedy decoding using Qwen2.5 14B as target and 0.5B as draft, and find that the speedup ratio of greedy decoding is much higher compared with the sampling experiments in Section 3.1.
+Further investigation suggests that this is a result of repetition hallucination in both target and draft models during long-form greedy generation.
+Table 5: Results of greedy decoding on MT-Bench. Greedy decoding leads to repetition hallucinations in both target and draft models in long-form generation, resulting in exaggerated speedup ratio.
+Model Methods Context Length
 128 256 512 1K 2K 4K 6K 8K
-Pythia (6.9B, ${}_{\text{(6.9B, 160M)}}$ Const. 1.10 1.30 1.50 1.66 1.82 1.20 1.11 1.04
-Heuristics 1.25 1.45 1.65 1.81 2.02 1.50 1.46 1.41
-SVIP 1.41 1.62 1.83 2.01 2.21 1.71 1.60 1.52
-Qwen2.5 (14B, ${}_{\text{(14B, 0.5B)}}$ Const. 1.05 1.08 1.15 1.29 1.44 1.54 1.60 1.67
+Qwen2.5 (14B, 0.5B) Const. 1.05 1.08 1.15 1.29 1.44 1.54 1.60 1.67
 Heuristics 1.04 1.06 1.13 1.32 1.54 1.72 1.85 1.97
 SVIP 1.30 1.34 1.42 1.57 1.74 1.87 1.98 2.10
-LLaMA-3 (70B, ${}_{\text{(70B, 8B)}}$ Const. 2.06 2.18 2.31 2.45 2.58 2.72 2.77 2.78
-Heuristics 2.26 2.46 2.73 3.07 3.48 3.90 4.15 4.26
-SVIP 2.31 2.56 2.86 3.21 3.59 4.00 4.23 4.33
-Table 6: Speedup on MT-Bench with different generation length using greedy decoding.
-Figure 8: The correlation between draft model entropy and lengths of accepted draft seqeunces in the greedy setting.
-In Table 6, we present the results of long-form generation with greedy decoding.
-Compared with Table 2, the speedup ratio of greedy decoding is much higher (even more than 4 times for LLaMA-3 after 4K context).
-This is due to the fact that these models tend to repeat themselves in greedy long-form generation, making it very easy for the draft models to predict the next tokens.
-In Figure 8, we also plot the relation between draft model entropy and accepted draft sequence lengths in this setting, which shows a much stronger correlation compared with the sampling setting in Figure 2.